Wheeling Manufacturing orders 8,000 units of graphite shafts for its production of golf clubs per week. The carrying costs of these shafts are $5 per unit per year and the fixed ordering cost is $300. What are the annual holding costs and annual ordering costs? Is ordering once a week too often or not often enough? How do you tell?
Answer #1

total carrying cost = 1/2 *no. of units per order* carrying cost per unit per year

= 1/2*8000*5 = 20000

total ordering cost = fixed ordering cost per order *no. of orders

= 300*52 =15600

to find out whether the ordering is too often or not we should calculate EOQ and compare the cost with the above scenario

EOQ = root of 2 AO/C or (2AO/C)-2

where, A = annual demand of the unit =8000*52 =416000

O = ordering cost per order = 300

C = carrying cost per unit per year = 5

hence, the EOQ= (2*416000*300/5)-2 = 7065.41

the total carrying cost =1/2*7065.41*5 =17663.5

total ordering cost = 300*no of orders

no.of orders = total demand/EOQ

=416000/7065.41 =58.88

hence, total ordering cost = 17663.5

hence we find that the total associated cost at EOQ is very lower hence the above scenario is too often and shall be decreased to 58 days